OpenStudy (anonymous):

An object having a mass of 2.0 kilograms falls from a height of 15 meters. What is its kinetic energy when it hits the ground?

OpenStudy (anonymous):

Use conservation of energy: mgh1 + 1/2 mv!^2 = mgh2 + 1/2 mv2^2 h2 = 0 v1= 0 So, the final kinetic energy: 1/2 mv2^2 = mgh1